OpenStudy (anonymous):

Solve the system by the substitution method. Show your work. 2y - x = 5 x2 + y2 - 25 = 0

OpenStudy (igreen):

Solve for 'x' in the first equation, can you do that?

OpenStudy (igreen):

2y - x = 5 -x = 5 - 2y x = 2y - 5 Now plug in 2y - 5 for 'x' in the 2nd equation and solve for 'y'.